### v1

# A Sampling-based Framework for Hypothesis Testing on Large Attributed Graphs [Scalable Data Science]

Abstract.
Hypothesis testing is a statistical method used to draw conclusions about populations from sample data, typically represented in tables. With the prevalence of graph representations in real-life applications, hypothesis testing in graphs is gaining importance. In this work, we formalize node, edge, and path hypotheses in attributed graphs. We develop a sampling-based hypothesis testing framework, which can accommodate existing hypothesis-agnostic graph sampling methods. To achieve accurate and efficient sampling, we then propose a Path-Hypothesis-Aware SamplEr, PHASE, an $m$-dimensional random walk that accounts for the paths specified in a hypothesis. We further optimize its time efficiency and propose $\text{PHASE}_{\text{opt}}$. Experiments on real datasets demonstrate the ability of our framework to leverage common graph sampling methods for hypothesis testing, and the superiority of hypothesis-aware sampling in terms of accuracy and time efficiency.
1. Introduction
Hypothesis testing finds widespread application in various domains such as marketing, healthcare, and social science (Newey and McFadden, 1994). Hypotheses are usually tested on representative samples since it is impractical or even impossible to extract data from entire populatations due to size, accessibility or cost. For instance, snowball sampling has been proven effective in accessing a hidden population like drug abusers through a chain-referral mechanism (Waters, 2015; Goodman, 1961). In this paper, we study hypothesis testing on graphs.
Graphs can be used to represent many real-world applications such as social, bibliographic and transportation networks where entities are nodes and edges reflect relationships between entities. Specifically, we work on Attributed Graphs that are graphs with multiple types of nodes and edges, both of which are labeled with attributes containing valuable information for graph analytics (Croft et al., 2011; Omidvar-Tehrani et al., 2019). In a bibliographic network like DBLP (see Figure 1a and Figure 1b for an example and an attributed path of length one, respectively), hypotheses expressed on nodes, edges, and paths unveil interesting aspects of collaboration behavior and research trends based on citation evolution over time. Specifically, a node hypothesis can be “the average citation of conference papers is greater than average”, and a path hypothesis can be “the minimum number of citations of conference papers written by Microsoft researchers is greater than average”. Similarly, on social platforms like Yelp, hypothesis testing can be used to check how different users review a product and study its adoption.
Figure 1. DBLP network schema and paths
Objectives and Challenges. Given an input attributed graph $\mathcal{G}$, our aim is to verify a hypothesis $H$ and return true or false. We hence seek to achieve two objectives: O1 - hypothesis testing should be as accurate as possible; O2 - execution time should be as small as possible. To the best of our knowledge, there is no existing literature addressing hypotheses on attributed graphs. A straightforward approach is to obtain a sampled graph $\mathcal{S}$ from $\mathcal{G}$ using hypothesis-agnostic graph sampling methods (Maiya and Berger-Wolf, 2010; Li et al., 2019; Adamic et al., 2001). However, the challenge is that depending on the sampling method and on the budget $B$, i.e., the size of $\mathcal{S}$, the hypothesis estimator, which computes aggregated values of nodes, edges, or paths specified in the hypothesis may return different values on $\mathcal{S}$ and $\mathcal{G}$. This raises the accuracy challenge behind O1. Additionally, as $B$ increases, the accuracy of the resulting hypotheses increases. However, the speed at which the accuracy converges highly depends on the sampling. This raises the efficiency challenge behind O2.
Contributions. We classify hypotheses on attributed graphs into three types: node, edge, and path hypotheses. We develop a sampling-based hypothesis testing framework, which accommodates common hypothesis-agnostic sampling methods, such as random node sampler (Stumpf et al., 2005), simple random walk (Gjoka et al., 2010), and non-backtracking random walk (Lee et al., 2012).
The lack of awareness of the input hypothesis in existing sampling methods may slow down the convergence speed of accuracy. Indeed, the sampled graph $\mathcal{S}$ should preserve as many relevant paths according to $H$ from $\mathcal{G}$ as possible for a path hypothesis. Therefore, to address objectives O1 and O2, we design PHASE, a Path-Hypothesis-Aware SamplEr. PHASE is aware of $H$ and preserves its nodes, edges, or paths. Consequently, the resulting $\mathcal{S}$ is more likely to provide relevant information for testing the hypothesis. PHASE employs $m\geq 1$ dependent random walks with two weight functions to ensure path-hypothesis-awareness. One function biases the seed selection toward the first node in the path hypothesis. The other biases the transition probability towards the nodes specified in the hypothesis. We further propose $\text{PHASE}_{\text{opt}}$ to improve the time efficiency of PHASE. $\text{PHASE}_{\text{opt}}$ reduces computation overhead through random sampling of neighbors. It also adopt a non-backtracking approach to prioritize the selection of relevant nodes, edges, and paths. We demonstrate both theoretically and empirically that the hypothesis estimator on $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
We conduct extensive experiments on three real-world datasets: MovieLens, DBLP, and Yelp. We aim to demonstrate the effectiveness of the optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE, and to compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance, accuracy, and running time.
We first observe that $\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE. Then, by comparing $\text{PHASE}_{\text{opt}}$ with 11 state-of-the-art hypothesis-agnostic samplers, we find it achieves the best accuracy given the sampling budget across various hypothesis types and datasets. $\text{PHASE}_{\text{opt}}$ also exhibits fast convergence of hypothesis estimators and robust accuracy performance especially for hypotheses with longer paths or limited relevant nodes, edges, or paths in $\mathcal{G}$. For significance, $\text{PHASE}_{\text{opt}}$ consistently provides the most precise and reliable estimate, achieving the highest hypothesis significance among all samplers. Moreover, $\text{PHASE}_{\text{opt}}$ achieves high accuracy in the shortest amount of time, making it usable in practice.
2. Definitions
2.1. Attributed Graphs
Definition 0 (Attributed Graph).
An attributed graph is a directed graph $\mathcal{G}=(\mathcal{V,E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges, represented by ordered pairs of nodes. It has a node type mapping function $\phi:\mathcal{V}\rightarrow\mathcal{T}$ and an edge type mapping function $\psi:\mathcal{E}\rightarrow\mathcal{R}$. Each node and edge has attributes.
Each node $v$ in $\mathcal{G}$ belongs to a specific node type $\phi(v)\in\mathcal{T}$, and each edge $e=(u,v)$ connecting nodes of type $\phi(u)$ to $\phi(v)$ belongs to a specific edge type $r=\psi(e)\in\mathcal{R}$. Once the relation $r$ exists, its inverse relation $r^{-1}$ naturally holds from $\phi(v)$ to $\phi(u)$. In this work, we assume each node in $\mathcal{G}$ has at least one incoming or outgoing edge. It implies the connectedness of the graph.
The DBLP Network in Figure 1a is an attributed graph containing four types of nodes and four types of edges. For example, paper nodes belong to the node type $\mathtt{paper}\in\mathcal{T}$ with five attributes, including title and year. The edge from paper to field of study (FOS) nodes is denoted by $\mathtt{WithDomain}\in\mathcal{R}$, with a weight indicating the paper’s relevance to a particular FOS. The reverse relationship $\mathtt{WithDomain}^{-1}\in\mathcal{R}$ also holds from FOS to paper nodes.
Definition 0 (Path).
A path $\mathcal{P}$ is defined as
$t_{1}\xrightarrow{r_{1}}\dots\xrightarrow{r_{l}}t_{l+1}$
where the node type $t_{i}$ and edge type $r_{i}$ can repeat; $l\geq 0$ is the length of $\mathcal{P}$. When $l=0$, $\mathcal{P}$ is a node and when $l=1$, it is an edge.
An attributed path is a path where each node has a list of attributes, referred to as a modifier. Figure 1b presents an example of a length-one attributed path, “conference papers written by Microsoft researchers”, and Figure 1c is an instance of that path.
Figure 2. The Sampling-based Hypothesis Testing Framework on attributed graphs.
2.2. Hypotheses on Attributed Graphs
We formally define path hypotheses on attributed graphs. Node and edge hypotheses are two special cases of path hypotheses.
Definition 0 (Path Hypothesis).
Given a path $\mathcal{P}=t_{1}\xrightarrow{r_{1}}t_{2}\xrightarrow{r_{2}}...\xrightarrow{r_{l}}t_{l+1}$ in $\mathcal{G}$, where $l\geq 1$, a path hypothesis is defined as:
[…]
In the DBLP network defined previously, co-authorship can be represented as a path:
$\mathcal{P}=\texttt{author}\xrightarrow{\mathtt{Authorship}^{-1}}\texttt{paper}\xrightarrow{\mathtt{Authorship}}\texttt{author}$
“The average citation of papers co-authored by Microsoft researchers is greater than 100” can be expressed as:
[…]
where $f_{\mathcal{P}}=\mathtt{paper}[\mathtt{citation}]$, ${\tt MSR}$ stands for microsoft research.
When $\mathcal{P}$ has length zero, a path hypothesis is reduced to a node hypothesis. An example of node hypothesis in the DBLP dataset can be “the average number of citations for conference papers is larger than 50” expressed as:
[…]
When $\mathcal{P}$ has length one, we refer it as an edge hypothesis. For instance, the edge hypothesis “the average $\mathtt{FOS\_weight}$ of conference papers on data mining is larger than 0.5” can be expressed as:
[…]
C (resp. DM) stands for conference papers (resp. DM data mining).
2.3. Problem Statement and Challenges
Given an attributed graph $\mathcal{G}$, our aim is to verify a node, edge, or path hypothesis $H$ and return true or false.
To address our problem, we formulate two objectives: O1 the hypothesis testing should be as accurate as possible; O2 the overall execution time should be as small as possible.
When testing a hypothesis such as “the average number of citations for conference papers is larger than 50”, a common approach includes collecting representative sample data and computing an aggregate value. In this case, the aggregate value corresponds to the average number of citations. However, directly collecting all relevant, i.e., satisfying the modifiers in the hypothesis, nodes, edges, or paths from $\mathcal{G}$ and computing an aggregate value can be time-consuming. Therefore, we adopt the commonly used approach of sampling the graph $\mathcal{G}$ into a smaller representative graph $\mathcal{S}$. This assumes a sampling budget $B$ that reflects the maximum size of $\mathcal{S}$. For the sake of simplicity, unless stated otherwise, we assume sampling an edge or a node requires the same unitary cost $1$.
A hypothesis estimator returns the aggregated value of nodes, edges, or paths in a hypothesis. The challenge of addressing O1 is that depending on the sampling method and the budget $B$, the hypothesis estimator obtained from $\mathcal{S}$ may differ from that of $\mathcal{G}$. We propose to first examine the applicability of existing sampling methods for graphs. Our observation is that the accuracy of those methods increases slowly as $B$ increases, resulting in longer times to achieve satisfactory accuracy. We conjecture that this is attributed to the lack of awareness of the underlying hypothesis during sampling. For instance, $\mathcal{S}$ should preserve as many relevant paths from $\mathcal{G}$ as possible for a path hypothesis. Hence, to address objective O2, we seek to design a sampling method that is aware of $H$ and preserves relevant nodes, edges, or paths in $H$. We further aim to demonstrate both theoretically and empirically that using our sampling, the hypothesis estimator on resulting $\mathcal{S}$ converges to its corresponding value in $\mathcal{G}$ as $B$ increases.
3. Sampling-based Hypothesis Testing
We propose a sampling-based hypothesis testing framework that supports both hypothesis-agnostic and hypothesis-aware graph samplers. After reviewing existing hypothesis-agnostic methods, we introduce our hypothesis-aware sampler, PHASE, in Section 3.2.1. It addresses objective O1 by ensuring convergence of hypothesis estimators, and objective O2 by incorporating hypothesis awareness into the sampling. Additionally, we implement two optimizations, $\text{PHASE}_{\text{opt}}$, to reduce the running time of PHASE in Section 3.2.2.
3.1. Hypothesis-Agnostic Samplers
Figure2 illustrates the sampling-based hypothesis testing framework we propose to test node, edge, and path hypotheses on attributed graphs. Based on whether the graph sampling method is hypothesis-agnostic or hypothesis-aware, there are two workflows indicated by dashed and dash-dot-dash arrows, respectively. Their distinction lies in where the hypothesis $H$ is introduced. In the hypothesis-aware workflow, $H$ is necessary during the sampling step, while in the hypothesis-agnostic workflow, $H$ is directly introduced before hypothesis testing.
Following dashed arrows from the left, the first procedure is to sample a graph $\mathcal{S}$ using a hypothesis-agnostic sampler, which requires $\mathcal{G}$ and $B$ only. The existing hypothesis-agnostic samplers can be categorised into three categories: node samplers that choose $B$ nodes from $\mathcal{G}$(Stumpf et al., 2005; Adamic et al., 2001), edge samplers that choose $B$ edges from $\mathcal{G}$(Krishnamurthy et al., 2005), and random walk based samplers that pick edges and nodes through the process of random walks(Gjoka et al., 2010; Lee et al., 2012; Leskovec and Faloutsos, 2006; Li et al., 2019). The second procedure is to conduct the hypothesis testing using $\mathcal{S}$ and $H$. Specifically, we extract all relevant nodes, edges, or paths from $\mathcal{S}$ and compute the required aggregated values for hypothesis testing. Finally, the acceptance or rejection result, p-value, and confidence interval are returned.
3.2. Hypothesis-Aware Samplers
3.2.1. PHASE Algorithm
We introduce PHASE, our Path-Hypothesis-Aware SamplEr (Algorithm1). PHASE aims to preserve nodes, edges, or paths specified in a hypothesis $H$. It can be integrated with any random walk based sampler that maintains the path-preserving property. To ease understanding, we illustrate it with FrontierS(Ribeiro and Towsley, 2010). FrontierS picks a node from $m$ random seeds based on degree-proportional probability and performs a random walk by uniformly selecting a neighbor to proceed. It repeats the process until reaching the sampling budget $B$.
Algorithm 1 PHASE
1:Input: $\mathcal{G}=(\mathcal{V,E})$, $B$, $H$
2:Output: a sampled graph $\mathcal{S}$
3:Initialize $L=(v_{1},v_{2},\ldots,v_{m})$ with $m$ randomly chosen nodes
4:$L_{w}=\text{AssignWeight}(L,H)$triangleright Assign $w_{h}$ to $x_{1}$ nodes and $w_{l}$ to other nodes.
5:$\mathcal{V}_{\mathcal{S}}=\{\}$, $\mathcal{E}_{\mathcal{S}}=\{\}$
6:while $B>m$ do
7:    Normalize $L_{w}$
8:    Select $v\in L$ with probability $L_{w}$
9:    $N\leftarrow N[v]$triangleright $N[v]$ is the set of neighbors of $v$.
10:    $N_{w}=\text{AssignWeight}(N,H)$triangleright Assign weights to nodes in $N$ according to the transition probability matrices.
11:    Select $u\in N$ with probability $N_{w}$
12:    $\mathcal{V}_{\mathcal{S}}$.update($v,u$)
13:    Replace $v$ by $u$ in $L$ and update $L_{w}$
14:    $B=B-1$
15:end while
16:return $\mathcal{S}=\{\mathcal{V}_{\mathcal{S}},\mathcal{E}_{\mathcal{S}}\}$
PHASE takes $\mathcal{G}$, $B$, and $H$ as inputs. For a node, edge, or path hypothesis, $H$ contains $\mathcal{P}$ with lengths zero, one, or more than one, respectively. It initializes $m$ seed nodes randomly in a list $L$ (Line 1), which increases the probability of selecting relevant nodes, edges, and paths from $\mathcal{G}$. Each seed is assigned a weight in $L_{w}$ (Line 2) to guide the selection of seed nodes for random walks later. These weights are determined based on heuristics, with nodes satisfying the first node modifier on $\mathcal{P}$ (denoted by $x_{1}$) assigned a higher weight $w_{h}$ and all other nodes assigned a lower weight $w_{l}$.
Lines 4-14 describe an iterative random walk to select nodes until $B$ is reached. During each iteration, the algorithm chooses a node $v$ from $L$ with probability $L_{w}$ (Line 6). Then, it selects one of $v$’s neighbors $u$ using a weighted selection process (Lines 7-9). Specifically, in Line 9, we adjust the transition probability to bias the random walks towards nodes, edges, or paths that satisfy $\mathcal{P}$ in $H$. The transition probability matrices for node, edge, and path hypotheses are depicted in Figures 3a, 3b, and 3c, respectively.
After adding nodes $u$ and $v$ to $\mathcal{V}_{S}$ (Line 10), the algorithm updates $L$ by replacing $v$ with $u$ (Lines 11) and decreases the sampling budget by one (Lines 12). This iteration continues until $B>m$. The resulting sampled graph $\mathcal{S}$ is the induced subgraph from $\mathcal{V}_{S}$.
$x_{1}$$y$$x_{1}$$w_{h}$$w_{l}$$y$$w_{h}$$w_{l}$(a)$x_{1}$$x_{2}$$y$$x_{1}$$w_{l}$$w_{h}$$w_{l}$$x_{2}$$w_{h}$$w_{l}$$w_{l}$$y$$w_{h}$$w_{l}$$w_{l}$(b)$x_{1}$$x_{2}$$x_{3}$$y$$x_{1},x_{1}$$w_{l}$$w_{h}$$w_{l}$$w_{l}$$x_{1},x_{2}$$w_{l}$$w_{l}$$w_{h}$$w_{l}$$x_{1},x_{3}$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$x_{1},y$$w_{h}$$w_{l}$$w_{l}$$w_{l}$$\vdots$(c)
Figure 3. Transition probability matrices for node (a), edge (b), and path ($l=2$) (c) hypotheses, where $w_{h}\geq w_{l}>0$ denote transition probabilities. $x_{i}$ represents nodes in $\mathcal{G}$ satisfying the $i$-th node modifier on $\mathcal{P}$, while $y$ represents nodes not satisfying any node modifier on $\mathcal{P}$. (a) and (b) involve 1st-order random walks, whereas (c) involves 2nd-order random walks because the probability of selecting the next node depends on both the current and previous nodes.
3.2.2. $\text{PHASE}_{\text{opt}}$ Algorithm
We further propose optimizations to reduce execution time. Algorithm 2 introduces two lines that replace line 7 in Algorithm 1. Firstly, in densely connected graphs, the computation of neighbor weights in step 8 of Algorithm 1 can be time-consuming due to the potentially large number of neighbors. To mitigate this, we randomly sample a subset of $\min\{|N|,n\}$ neighbors as candidates for selection, where $n$ is a parameter. Secondly, we adopt a non-backtracking approach by avoiding the selection of visited nodes during the sampling process(Lee et al., 2012). This strategy prevents cycles and unnecessary node visits.
Algorithm 2 $\text{PHASE}_{\text{opt}}$
(showing only 2 lines that replace line 7 in Algorithm 1
1:$N^{\prime}\leftarrow N[v]-\mathcal{V_{S}}$ (Optim 2)
2:$N\leftarrow$ Select $\text{min}\{|N^{\prime}|,n\}\text{ nodes randomly from }N^{\prime}$ (Optim 1)
Complexity. On average, each node has $2|\mathcal{E}|/|\mathcal{V}|$ neighbors to be examined by PHASE, resulting in a time complexity of $O(B\times 2|\mathcal{E}|/|\mathcal{V}|)$. On the other hand, by constraining the number of neighbors, $\text{PHASE}_{\text{opt}}$ achieves a time complexity of $O(B)$.
3.2.3. Convergence of Hypothesis Estimators
Let us demonstrate the convergence of hypothesis estimators using our proposed sampling algorithms (Challenge 1).
By Lemma 5.1 from(Ribeiro and Towsley, 2010), FrontierS is equivalent to the sampling process of a SRW over $\mathcal{G}^{m}=(\mathcal{V}^{m},\mathcal{E}_{m})$, where $\mathcal{G}^{m}$ is the $m$-th Cartesian power of $\mathcal{G}$ and
$\mathcal{V}^{m}=\{(v_{1},\ldots,v_{m})|v_{1}\in\mathcal{V}\wedge\dots\wedge v_{m}\in\mathcal{V}\}$
is the $m$-th Cartesian power of $\mathcal{V}$. $\forall\boldsymbol{u,v}\in\mathcal{V}^{m}$, $(\boldsymbol{u,v})\in\mathcal{E}_{m}$ if there exists $i$ such that $(u_{i},v_{i})\in\mathcal{E}$ and for all $j\neq i$, $u_{j}=v_{j}$. Given $\mathcal{G}$ is a directed, connected, and non-bipartite graph, in the steady state, the sampled edges of FrontierS form a stationary sequence and satisfy the following Strong Law of Large Numbers (SLLN), the proof of which can be found in (Ribeiro and Towsley, 2010).
Theorem 1 (SLLN).
For any function $f$, where
$\sum_{(u,v)\in\mathcal{E}}|f(u,v)|<\infty$,
$\lim_{B\to\infty}\frac{1}{B}\sum^{B}_{i=1}f(u_{i},v_{i})\to\frac{1}{|E|}\sum_{\forall(u,v)\in\mathcal{E}}f(u,v)$
almost surely, i.e. the event occurs with probability one.
For a path hypothesis with length $l$, it involves $(l-1)$-th order random walk, which is equivalent to a SRW over $\mathcal{G}^{l-1}=(\mathcal{V}^{l-1},\mathcal{E}_{l-1})$(Lovász, 1993). Hence, PHASE’s random walks can be viewed as a SRW over $\mathcal{G}^{l-1+m}=(\mathcal{V}^{l-1+m},\mathcal{E}_{l-1+m})$. As PHASE ensures that the resulting random walks over $\mathcal{G}^{l-1+m}$ maintain both irreducibility and aperiodicity, it preserve stationarity and adhere to the SLLN. $\text{PHASE}_{\text{opt}}$ also follows the SLLN because the stationary distribution of the NBRW remains the same as that of the SRW.
By Theorem1, we can construct hypothesis estimators for node, edge, and path hypotheses and prove their convergence. Here, we focus on scenarios where agg represents an average function. The hypothesis estimators for other aggregation functions, such as maximum and minimum, can be derived in a similar manner.
A path hypothesis ($l\geq 2$): $\text{avg}(f_{\mathcal{P}}|M_{t_{i}},\forall t_{i}\text{ on }\mathcal{P})$ has a primary subject $f_{\mathcal{P}}$. Let $\mathcal{P}^{*}$ be all relevant paths in $\mathcal{G}$, and $\mathcal{P_{G}}$ be all paths with the same length as $\mathcal{P}$ in $\mathcal{G}$. The mean value of the path hypothesis, $\theta_{path}$, is
(1) $\theta_{path}=\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
where $T(\mathcal{P})=f_{\mathcal{P}}\times\mathbbm{1}_{M_{t_{i}}\subseteq\mathcal{L}_{\phi(t_{i})}\forall t_{i}\text{ on }\mathcal{P}}$. Replacing $\mathcal{G}$ with the sampled graph $\mathcal{S}$, the estimator for $\theta_{path}$ is
(2) $\hat{\theta}_{path}=\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})$
When $B$ goes to infinity, $\mathcal{S}$ converges to $\mathcal{G}$. By Theorem1, we have
(3) $\lim_{B\to\infty}\frac{1}{|\mathcal{P}^{*}\cap\mathcal{P_{S}}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{S}}}T(\mathcal{P})\to\frac{1}{|\mathcal{P}^{*}|}\sum_{\forall\mathcal{P}\in\mathcal{P_{G}}}T(\mathcal{P})$
almost surely. Hence, $\hat{\theta}_{path}$ is an asymptotically unbiased estimator of $\theta_{path}$. Hypothesis estimators for node and edge hypotheses can be derived similarly.
4. Experiments
The goal of our experiments is twofold: 1) test the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE (Section 4.3), and 2) compare hypothesis-agnostic and hypothesis-aware samplers in terms of test significance (Section 4.4), accuracy (Section 4.5), and time (Section 4.6). The source code is publicly available  [^1] https://github.com/Carrieww/GraphHT.
4.1. Experimental Setup
Table 1. Statistics of datasets
Dataset #(Nodes) #(Edges) Density #(Node Types) #(Edge Types)
MovieLens 9,705 996,656 1.06e-02
2
1
DBLP 1,623,013 11,040,170 4.19e-06
4
4
Yelp 2,136,118 6,743,879 1.48e-06
2
1
Datasets. We used three datasets (Harper and Konstan, 2016; Tang et al., 2008), extracted from real attributed networks: MovieLens, DBLP, and Yelp. Table 1 shows the statistics of the datasets.
Samplers We compare $\text{PHASE}_{\text{opt}}$ with eleven existing samplers:
  • Node samplers: Random Node Sampler (RNS), Degree-Based Sampler (DBS)
  • Edge samplers: Random Edge Sampler (RES)
  • Random walk based samplers: Simple Random Walk (SRW), Frontier Sampler (FrontierS), Non-Backtracking Random Walk (NBRW), Random Walk with Restarter (RWR), Metropolis-Hastings Random Walk (MHRW), Snaw Ball Sampler (SBS), Forest Fire Sampler (FFS), Shortest Path Sampler (ShortestPathS)
Hypotheses. For each dataset (DB) and hypothesis type (N for node hypotheses), we provide three example hypotheses (labeled as DB-N1) in experiments. For example, on Yelp, we explore an edge hypothesis testing if “fast food average ratings exceed 4”. In DBLP, we examine a path hypothesis to see if “papers co-authored by Chinese institutes on data mining have over 50 citations”. To assess path length effects, we further test two path hypotheses with lengths of three and four in DBLP.
Parameter Choice. For PHASE and $\text{PHASE}_{\text{opt}}$, We set $m=50$ to maintain the path-preserving capability of multi-dimensional random walks, and $n=30$ to strike a balance between accuracy and time efficiency. Additionally, we choose $w_{h}=10$ and $w_{l}=0.1$ to regulate the bias towards relevant nodes, edges, and paths. The sampling budget $B$ is maintained as a proportion of the total number of nodes in $\mathcal{G}$ for all samplers. As for existing hypothesis-agnostic samplers, we use the best parameters in their respective settings. We report an average of 30 runs for every evaluation measure.
4.2. Evaluation Measures
Accuracy. Accuracy at a specific sampling proportion reflects the effectiveness of a sampling method. The accuracy below measures the number of matched hypothesis testing results on $\mathcal{G}$ and $\mathcal{S}$:
$\text{Accuracy}=\frac{1}{k}\sum_{k}\mathbbm{1}_{H(\mathcal{G})==H(\mathcal{S})}$
where $H(\mathcal{G})$ and $H(\mathcal{S})$ return 0 (false) or 1 (true).
Time. We measure the total execution time, including both the sampling time and the time taken to extract relevant information from $\mathcal{S}$ for hypothesis testing.
Table 2. Average execution time (sec) of PHASE and $\text{PHASE}_{\text{opt}}$ for DBLP node, edge, and path hypotheses.
Node Edge Path
PHASE 115.66 539.15 441.3
$\text{PHASE}_{\text{opt}}$ 5.56 8.76 5.91
Table 3. The accuracy of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset
Hypothesis Type
Sampling Proportion (%)
$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRW ShortestPathS FrontierS FFS SBS
Node
1
0.9 0.89 0.83 0.86 0.88 0.89 0.87 0.89 0.86 0.84 0.83 0.56
Edge
2.5
0.98 0.68 0.77 0.98 0.99 1 0.99 0.98 0.99 0.73 0.93 0.91
MovieLens
Path
5
0.99 0.1 0.88 0.83 0.82 0.89 0.91 0.98 0.95 0.38 0.85 0.62
Node
0.2
0.96 0.48 0.87 0.93 0.92 0.91 0.9 0.94 0.92 0.92 0.94 0.88
Edge
0.2
0.76 0.48 0 0.71 0.73 0.69 0.7 0.29 0.69 0.63 0.7 0.42
DBLP
Path
0.2
0.89 0 0 0 0.26 0.3 0.32 0.18 0.33 0.043 0.3 0.12
Node
0.1
0.99 0.65 0.77 0.61 0.69 0.69 0.69 0.77 0.7 0.64 0.66 0.48
Edge
1
1 0.73 0.54 0.76 0.79 0.91 0.87 0.84 0.76 0.77 0.79 0.71
Yelp
Path
1
0.99 0.11 0.05 0.79 0.78 0.78 0.72 0.8 0.99 0.42 0.67 0.42
Table 4. The execution time (sec) of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ on three datasets and three types of hypothesis.
Dataset
Hypothesis Type
Sampling Proportion (%)
$\text{PHASE}_{\text{opt}}$ RES RNS DBS SRW NBRW RWR MHRWS ShortestPathS FrontierS FFS SBS
MovieLens
Node
1
0.083 0.99 0.023 0.077 0.057 0.06 0.05 0.06 0.063 0.083 0.067 0.047
Edge
2.5
0.45 0.99 0.11 0.36 0.31 0.36 0.33 0.29 0.26 0.33 0.35 0.31
Path
5
4.92 1.03 0.34 4.32 4.53 4.55 4.76 3.10 0.95 0.38 3.87 3.08
DBLP
Node
0.2
5.56 18.70 0.55 6.98 7.98 8.22 9.67 1.66 31.32 10.48 5.73 3.03
Edge
0.2
8.76 22.57 0.90 10.07 14.76 12.89 12.32 3.53 33.91 14.46 8.87 5.27
Path
0.2
5.44 19.61 0.71 6.87 8.24 8.39 8.13 2.17 31.13 9.37 5.33 3.01
Yelp
Node
0.1
2.56 13.84 1.02 6.81 1.19 0.96 1.16 1.12 1.84 1.68 1.02 6.58
Edge
1
19.42 16.88 2.01 13.61 8.58 8.97 9.17 6.93 30.23 10.22 10.27 6.03
Path
1
56.97 14.35 1.53 23.48 15.97 16.61 19.03 7.35 37.81 9.95 25.61 19.94
4.3. PHASE vs $\text{PHASE}_{\text{opt}}$
We evaluate the effectiveness of optimizations in $\text{PHASE}_{\text{opt}}$ compared to PHASE using the DBLP dataset, as shown in Table 2. $\text{PHASE}_{\text{opt}}$ is at least 20 times faster than PHASE while maintaining comparable accuracy (less than 5% accuracy reduction). Consequently, we exclusively use $\text{PHASE}_{\text{opt}}$ for subsequent experiments.
4.4. Significance
(a) P-value Plot
(b) CI Plot
Figure 4. DBLP p-value and CI plot for the hypothesis DB-P1
To examine statistical significance and estimation precision, we analyze the trend of p-values and confidence intervals (CI) with increasing $B$. We illustrate this trend using a path hypothesis on DBLP as all other hypotheses show similar trends. As shown in Figure 3(a), the p-value of $\text{PHASE}_{\text{opt}}$ consistently stays below the significance level (e.g., 0.05) and decreases gradually. This trend indicates stronger evidence against the null hypothesis, showing higher statistical significance as $B$ increases.
Figure 3(b) shows narrowing CI with increasing $B$, indicating more precise and reliable estimations of the hypothesis estimator. $\text{PHASE}_{\text{opt}}$ consistently achieves the narrowest bound at any $B$, indicating the most precise and reliable estimate among all samplers.
(a) Accuracy ($l=3$)
(b) Accuracy ($l=4$)
Figure 5. Accuracy versus sampling proportion plots for DBLP path hypotheses with length three (a) and four (b).
4.5. Accuracy
Table 3 presents the accuracy performance of eleven existing samplers and $\text{PHASE}_{\text{opt}}$ at a specific sampling proportion, shown in column three. We select this proportion to balance stability and discriminative power, ensuring the accuracy is sufficiently stabilized without saturating across all samplers. For each dataset and type of hypothesis, the accuracy is the average accuracy of three hypothesis examples, each derived from an average of 30 runs. The highest accuracy in each row is highlighted in bold.
In Table 3, $\text{PHASE}_{\text{opt}}$ demonstrates robust performance in all cases, except for MovieLens edge hypotheses, in where $\text{PHASE}_{\text{opt}}$ slightly trails behind NBRW. Comparing among columns, some random walk based samplers, namely SRW, RWR, and ShortestPathS, exhibit good accuracy performance for edge and path hypothesis. On the contrary, RES, RNS, and DBS show ineffectiveness in handling path hypotheses. These observations align with their sampling mechanisms. These node and edge samplers can hardly preserve the path information from $\mathcal{G}$ in the sampled graph. Additionally, comparing $\text{PHASE}_{\text{opt}}$ and FrontierS, $\text{PHASE}_{\text{opt}}$’s superior performance indicates the efficacy of the two weight functions in enhancing the sampler’s hypothesis-awareness to sample relevant nodes, edges, and paths.
(a) Time $(l=3)$
(b) Time ($l=4$)
Figure 6. Time versus accuracy plots for DBLP path hypotheses with length three (a) and four (b).
(a) DB-N1
(b) DB-N2
(c) DB-N3
(d) DB-N3
(e) DB-E1
(f) DB-E2
(g) DB-E3
(h) DB-E3
(i) DB-P1
(j) DB-P2
(k) DB-P3
(l) DB-P3
Figure 7. Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the DBLP dataset.
We establish an accuracy ranking for all samplers by averaging their accuracy performance per column in Table 3 and select the top five: $\text{PHASE}_{\text{opt}}$, NBRW, ShortestPathS, RWR, and SRW. These samplers’ accuracy convergence on individual hypotheses is compared in Figures 7 and 8 for DBLP and Yelp datasets, respectively. The figure for MovieLens is excluded due to space constraints and the similarity of results. In these figures, a-c, e-g, and i-k display the accuracy performance of three node, three edge, and three path hypotheses, respectively. Hypotheses of the same type are structured to have progressively fewer relevant nodes, edges, or paths in $\mathcal{G}$. The x-axis is truncated to clearly show the convergence.
In Figures 7 and 8, $\text{PHASE}_{\text{opt}}$ consistently outperforms other methods across all sampling proportions for most hypotheses. For hypotheses of the same type, $\text{PHASE}_{\text{opt}}$, with hypothesis-awareness, consistently demonstrates robust accuracy performance for the hypotheses with fewer relevant nodes, edges, or paths in $\mathcal{G}$. This is particularly evident for DBLP edge and path hypotheses. When relevant nodes, edges, and paths in $\mathcal{G}$ are abundant, conventional sampling methods may have sampled a sufficient number of relevant nodes, edges, or paths to achieve high accuracy, which can diminish the distinctive advantage of $\text{PHASE}_{\text{opt}}$.
Delving into Figure 7, $\text{PHASE}_{\text{opt}}$ dominates in node and path hypotheses for the DBLP dataset. However, ShortestPathS slightly outperforms $\text{PHASE}_{\text{opt}}$ in edge hypotheses one and two, possibly due to the high betweenness centrality of these edges, implying that many shortest paths traverse those edges. Moving to Figure 8, $\text{PHASE}_{\text{opt}}$ demonstrates clear superiority across all types of hypotheses for the Yelp dataset. Nonetheless, ShortestPathS occasionally exhibits competitive performance compared to $\text{PHASE}_{\text{opt}}$ in path hypotheses, which may also be attributed to the betweenness centrality of those paths in $\mathcal{G}$.
Figure 5 shows how path length affects accuracy using the DBLP dataset. $\text{PHASE}_{\text{opt}}$ enables faster convergence of accuracy with longer path lengths, outperforming other samplers.
4.6. Scalability
(a) YP-N1
(b) YP-N2
(c) YP-N3
(d) YP-N3
(e) YP-E1
(f) YP-E2
(g) YP-E3
(h) YP-E3
(i) YP-P1
(j) YP-P2
(k) YP-P3
(l) YP-P3
Figure 8. Comparison of the top five sampling methods for accuracy and time (d, h, l) efficiency across three node (a, b, c), three edge (e, f, g), and three path hypotheses (i, j, k) for the Yelp dataset.
In Table 4, samplers’ execution time increases as the dataset becomes larger. Among them, RNS generally exhibits the shortest execution time as it uniformly samples nodes randomly. The performance of $\text{PHASE}_{\text{opt}}$ varies across datasets. For the Yelp dataset, it ranks among the highest, particularly for the path hypothesis. This can be attributed to the limited node types in Yelp, resulting in longer extraction times for condensed relevant paths from $\mathcal{S}$. For the DBLP dataset, $\text{PHASE}_{\text{opt}}$ ranks among the lowest five. However, instead of considering execution time alone, we should consider both execution time and accuracy performance to evaluate a sampler’s efficiency and effectiveness.
For the top five samplers ranked by their accuracy performance, we plot their accuracy versus execution time in subfigures d, h, and l of Figures 7 and 8. Due to space limitations, only the third hypothesis, which consists of the least number of relevant nodes, edges, or paths in the input graph, of each type is included. First of all, we observe that the curve of $\text{PHASE}_{\text{opt}}$ is always at the bottom among others. This means that $\text{PHASE}_{\text{opt}}$ consistently achieves high accuracy in the shortest amount of time across all plots. Additionally, its execution time does not exhibit exponential growth compared to other existing hypothesis-agnostic samplers. This can be attributed to $\text{PHASE}_{\text{opt}}$’s hypothesis-awareness, which allows it to efficiently sample relevant nodes, edges, and paths from the input graph to verify the hypothesis.
The impact of path length on accuracy and execution time is illustrated in Figure 6. As the path length increases, $\text{PHASE}_{\text{opt}}$ consistently achieves the highest accuracy within the shortest execution time.
5. Related Work
There has been a lot of research interest on hypothesis testing (Hunter et al., 2008; Gibbs et al., 2022; Arias-Castro and Verzelen, 2014; Xia and Li, 2017; Croft et al., 2011; Ghoshdastidar and von Luxburg, 2018; Omidvar-Tehrani et al., 2019), proposing sampling techniques on graphs (Leskovec and Faloutsos, 2006; Lee et al., 2012; Li et al., 2019). In this section we summarize the most representative works.
Hypothesis Testing on Graphs. Hypotheses on graphs can be categorized by their object of interest (Zhao et al., 2011; Arias-Castro and Verzelen, 2014; Bickel and Sarkar, 2013). Tang et al. and Ghoshdastidar et al. (Tang et al., 2017; Ghoshdastidar et al., 2017, 2020) extend the one-sample problem into two-sample to test whether two groups of random graphs are obtained by the same generative model or with the same graph distribution. In (Xia and Li, 2017), nodes are objects of interest, and one-sample hypothesis testing is used to detect conditional dependence between nodes in brain connectivity networks.
Our goal is more related to works when the object of interest is nodes. However, there are two differences between our work and the existing ones. First, we enable more expressive hypotheses, including edge and path hypothesis, on graphs. Second, we leverage graph sampling methods on graph hypothesis testing.
Sampling Methods on Graphs. Most sampling methods are designed to sample a representative subgraph to accurately estimate graph properties (Leskovec and Faloutsos, 2006; Stutzbach et al., 2009; Maiya and Berger-Wolf, 2010; Hübler et al., 2008). Leskovec and Faloutsos (Leskovec and Faloutsos, 2006) are the first to study this problem in real-world networks (Leskovec and Faloutsos, 2006) by proposing sampling techniques to maintain degree distribution, clustering coefficient and distribution of component sizes. Hübler et al. (Hübler et al., 2008) propose the Metropolis-Hastings sampling method. Later, Maiya and Berger-Wolf (Maiya and Berger-Wolf, 2010) propose a community representativeness sample and design a community structure expansion sampler. Many sampling methods are proposed to improve the efficiency and convergence rate of simple random walk (Lee et al., 2012; Li et al., 2019). Besides representative sampling, other sampling methods are designed for specific tasks, such as graph compression (Besta et al., 2019), community detection(Maiya and Berger-Wolf, 2010), and graph visualization (Rafiei and Curial, 2005; Wu et al., 2017). However, none of them is designed for hypothesis testing on attributed graphs.
6. Conclusion
In this paper, we developed a framework for hypothesis testing on large attributed graphs. Our framework accommodates eleven common hypothesis-agnostic samplers as well as new hypothesis-aware samplers. We proposed dedicated optimizations to speed up sampling while achieving high test significance and accuracy. We demonstrated theoretically and empirically that the convergence rate of our hypothesis-aware sampling.
Our work opens several new directions in the area of hypothesis sampling on graphs. The first direction is to examine additional optimizations that make use of domain-specific information on the input graph. The second direction is to handle more expressive hypotheses and specify two-sample and multiple sample scenarios.